# Density Fluctuations in the Intracluster Medium: An Attempt to Constrain Viscosity with Cosmological Simulations

### v1

@@ Abstract
The impact of viscosity in the Intracluster Medium (ICM) is still an open question in astrophysics. To address this problem, we have run a set of cosmological simulations of three galaxy clusters with a mass larger than $M_{\mathrm{Vir}}>10^{15}$ at $z=0$ using the SPMHD-code OpenGadget3. We aim to quantify the influence of viscosity and constrain its value in the ICM. Our results show significant morphological differences at small scales, temperature variations, and density fluctuations induced by viscosity. We observe a suppression of instabilities at small scales, resulting in a more filamentary structure and a larger amount of small structures due to the lack of mixing with the medium. The conversion from kinetic to internal energy leads to an increase of the virial temperature of the cluster of $\sim$5- 10%, while the denser regions remain cold. The amplitude of density fluctuations is found to increase with viscosity, as well as the velocity fluctuations. However, comparison with observational data indicates consistency with observed density fluctuations, challenging the direct constraint of viscosity solely through density fluctuations. Furthermore, the ratio of density to velocity fluctuations remains close to 1 regardless of the amount of viscosity, in agreement with the theoretical expectations. Our results show for the first time in a cosmological simulation of a galaxy cluster the effect of viscosity in the ICM, a study that is currently missing in the literature.
@@ 1 Introduction
The Intracluster Medium (ICM) is the most dominant baryonic visible component in galaxy clusters, filling the gravitational potential with hot ($T\sim 10^{7}-10^{8}$ K) ionised, X-ray emitting gas. This gas is continuously perturbed by galaxy motions (e.g. Faltenbacher et al., 2005), mergers (e.g. ZuHone, 2011; Iapichino et al., 2017), AGN outflows (e.g. O’Neill et al., 2009; Gaspari, 2015) and accretion of gas along filaments (e.g. Kravtsov & Borgani, 2012; Vallés-Pérez et al., 2021a). These energetic processes inject energy at large scales (low wavenumber $k$), which decay towards higher $k$ modes in a Kolmogorov-like cascade, introducing turbulence on a wide range of scales (e.g. Kolmogorov, 1941, 1962). The energy is then dissipated into heat, affecting small scale processes such as cosmic ray re-acceleration (e.g. Fujita et al., 2003; Brunetti & Lazarian, 2007), star formation (e.g. Federrath, 2016) or magnetic field amplification (e.g. Kazantsev, 1968; Kulsrud & Anderson, 1992). These complex mechanisms over cosmological periods of time produce very complicated scenarios in which the gas properties are fundamental.
@@
The properties of these density perturbations do not only reflect the driving mechanisms that trigger them, but also depend on the microphysics of the ICM, specifically, for thermal conduction (Ruszkowski et al., 2011) and viscosity (Zhuravleva et al., 2019; Kunz et al., 2022). Zhuravleva et al. (2019) suggested to use observations of density fluctuations to constrain the amount of viscosity in galaxy clusters. The idea is to compare the amplitude of density fluctuations measured in observations against simulations with different amounts of viscosity. Using Coma observations and simulations of incompressible hydrodynamic turbulence carried out using Direct Numerical Simulations (DNS), Zhuravleva et al. (2019) concluded that the effective isotropic viscosity must be suppressed by at least a factor of $\sim$10 to $\sim$1000 with respect to the Spitzer value. This suppression could be the effect of magnetic fields, which reduce the effect of viscosity (ZuHone et al., 2015; Berlok et al., 2019; Squire et al., 2023), as well as an enhanced collision rate due to plasma instabilities (Schekochihin & Cowley, 2006; Kunz et al., 2014; Berlok et al., 2021).
@@
The physics of the ICM has been studied extensively using simulations of isolated systems with more idealised setups (e.g. ZuHone et al., 2009; Brüggen et al., 2012; Mohapatra et al., 2020; Zhang et al., 2020) and cosmological simulations (e.g. Dolag et al., 2005; Teyssier et al., 2011; Vazza et al., 2016; Pakmor et al., 2023; Steinwandel et al., 2023). Similarly, viscosity in the context of ICM has also been extensively studied with simulations, focusing on its effect on AGN-powered bubbles (Scannapieco & Brüggen, 2008; Dong & Stone, 2009), on the suppression of instabilities in cold fronts (Roediger & ZuHone, 2011; Suzuki et al., 2013; ZuHone et al., 2015) and on its impact on stripped galaxies (Roediger et al., 2015; Kraft et al., 2017).
@@
However, viscosity in the intracluster medium (ICM) has not been extensively studied using cosmological simulations. The only study in the literature, by Sijacki & Springel (2006), did not provide a detailed analysis of these effects. In our study, we conduct a set of cosmological simulations focusing on how viscosity affects massive clusters in a realistic scenario. We use different values of the isotropic Spitzer viscosity (Spitzer, 1962; Braginskii, 1965) and compare it with the non-viscous case. Our goal is to provide a general view of how the morphology of the cluster is affected by viscosity and how the virial temperature depends on the amount of viscosity. We also want to test the criteria to constrain viscosity suggested by Zhuravleva et al. (2019) and if this criterion holds in the context of the complex, realistic cosmological simulation. To this end, we compare our results with observations of density fluctuations from Heinrich et al. (2024). Additionally, we want to verify if the density fluctuations scale linearly with the velocity fluctuations, independently of the how viscous the ICM is. Although full Spitzer viscosity is a larger value of viscosity than the one expected, due to the different mechanisms that suppress the effect of viscosity, we want to show this extreme case for a better understanding of the ongoing mechanisms. In a follow up work, we will show additional consequences of a viscous ICM such as changes of the magnetic field amplification, effects in the merger history or the different energy distribution, together with the time evolution of gas properties.
@@ 2.1 Theoretical considerations
$\rho\frac{\mathrm{d}\mathbf{v}}{\mathrm{d}t}+\nabla P=-\rho\nabla\Phi-\nabla\cdot\mathbf{\sigma}+\frac{(\nabla\times\mathbf{B})\times\mathbf{B}}{4\pi},$ (2)
$\frac{\mathrm{d}E}{\mathrm{d}t}+\mathbf{v}\cdot\nabla P+(E+P)\nabla\cdot\mathbf{v}-\nabla\cdot\frac{\mathbf{B}(\mathbf{v}\cdot\mathbf{B})}{4\pi}=\\ -\rho\mathbf{v}\cdot\nabla\Phi-\nabla\cdot\mathbf{Q}-\nabla\cdot\left(\mathbf{\sigma}\cdot\mathbf{v}\right)$ (3)
@@
is the Lagrangian derivative. $\rho$ is the gas density, $\mathbf{v}$ the velocity of the fluid, $\mathbf{B}$ the magnetic field, $\mathbf{Q}$ is the thermal conduction and $\Phi$ is the gravitational potential. $E$ is the total energy per unit volume
@@
with an adiabatic index of $\gamma=5/3$ for monoatomic gases. $\mathbf{\sigma}$ is the viscous stress tensor, which defines the behaviour of Spitzer viscosity as
$\mathbf{\sigma}=\eta\,\sigma_{ij}=\eta\,\left(\frac{\partial v_{i}}{\partial x_{j}}+\frac{\partial v_{j}}{\partial x_{i}}-\frac{2}{3}\delta_{ij}\frac{\partial v_{k}}{\partial x_{k}}\right)\,,$ (8)
where we have dropped the bulk viscosity term, since it is related to the degree of freedom of rotation of molecules, being zero for monoatomic gases (see e.g Zeldovich & Raizer, 1967; Pitaevskii & Lifshitz, 1981). The viscous stress tensor has a component ($\sigma_{ij}$) multiplied by the shear viscosity coefficient $\eta$, defined as
@@
Assuming subsonic gas motions within the galaxy cluster ($v\ll c_{s}$), density perturbations are expected to be proportional to the one-component velocity at each scale $l=1/k$. At large scales, assuming local pressure balance ($\delta P/P\ll\delta\rho/\rho$) and an equation of state $s=P/\rho^{\gamma}$, the small density perturbations can be written as
$\left(\frac{\delta\rho}{\rho}\right)^{2}=\left(\frac{1}{\gamma}\frac{\delta s}{s}\right)^{2}\,,$ (10)
where $s$ is the specific entropy of the gas. The restoring buoyancy force over a gas element will create an oscillatory movement in the radial direction dependent on the scale. Considering this, we can rewrite equation 10 as
$\left(\frac{\delta\rho}{\rho}\right)^{2}=\left(\frac{v_{r}}{\gamma\omega H_{s}}\right)^{2}=\frac{H_{p}}{H_{s}}\frac{k^{2}}{k_{\perp}^{2}}\left(\frac{v_{r}}{c_{s}}\right)^{2}\,,$ (11)
where $v_{r}$ is the radial velocity, $\omega$ is the oscillation frequency, $k$ is the wavenumber of the oscillations, $k_{\perp}$ is its projection perpendicular to the radial direction, $H_{p}=(\mathrm{d}\ln P_{0}/\mathrm{d}r)^{-1}$ the pressure scale height and $H_{s}=(\mathrm{d}\ln s_{0}/\mathrm{d}r)^{-1}$ is the entropy scale height (see Zhuravleva et al. (2014) for a detailed derivation). Assuming a stratified atmosphere, the motions are dominated by the perpendicular component of the velocity ($v_{1\mathrm{D}}\sim v_{\perp}\sim(k_{r}/k_{\perp})v_{r}\gg v_{r}$) and we can write equation 11 as
$\left(\frac{\delta\rho}{\rho}\right)_{k}^{2}=\eta_{\rho}^{2}\left(\frac{v_{1\mathrm{D}}}{c_{s}}\right)_{k}^{2}=\eta_{\rho}^{2}\,\mathcal{M}_{\mathrm{1D},k}^{2}$ (12)
At small scales (i.e. smaller than the Ozmidov scale [^2] The Ozmidov scale (Ozmidov, 1992) is defined as the scale where the turbulent shear and the stratification terms are equal in magnitude.), the turbulent cascade becomes isotropic ($k_{r}\sim k_{\perp}$ and $v_{r}\sim v_{\perp}$) and the density fluctuations become a passive scalar field. Thus, its variance is conserved and can be written as
$\left(\frac{\delta\rho}{\rho}\right)_{k}^{2}\sim\frac{\varepsilon_{\rho}\tau_{k}}{c_{s}^{2}}\,,$ (13)
which describes a flux $\varepsilon_{\rho}$ over the cascade time $\tau_{k}$ at which ($\delta\rho/\rho$) is mixed, normalised by the soundspeed. The turbulent motions decay at a speed
$v_{k}^{2}\sim\varepsilon_{v}\tau_{k}\,,$ (14)
where $\varepsilon_{v}$ is the kinetic flux. Plugging equation 14 into equation 13, the density fluctuations can be written as
$\left(\frac{\delta\rho}{\rho}\right)_{k}^{2}\sim\frac{\varepsilon_{\rho}}{\varepsilon_{v}}\left(\frac{v}{c_{s}}\right)_{k}^{2}\,.$ (15)
$\varepsilon_{\rho}/\varepsilon_{v}$ depends on how much variance of the scalar field reaches the inertial range from the larger scales, which must be equal to equation 12 at the Ozmidov scale. For a more detailed derivation of equations 12 and 15, see Zhuravleva et al. (2014).
@@ 2.2 Viscosity saturation
A proper treatment of viscosity requires the implementation of a viscosity saturation to avoid unphysical results. This is produced when the velocity length scale at which the momentum propagates due to viscosity is smaller than the ion mean free path. This means that the viscous momentum transfer proceeds faster than the mean soundspeed of the medium, overestimating this momentum transfer (Sarazin, 1986).
To avoid this, we need to introduce a viscosity saturation that limits the momentum propagation and ensures a smooth transition from the non-saturated to the saturated state. We do this in such a way that the momentum transfer of the saturated viscosity propagates at a velocity comparable to the soundspeed of the medium. We define the velocity length scale as $l_{v}=2\,c_{s}/|\sigma_{ij}|$, where $|\sigma_{ij}|=\sqrt{\sigma_{ij}^{2}}=\sqrt{\mathrm{tr}(\sigma_{ij}^{2})}$ is the strength of the tensor component of the viscous stress tensor. We now introduce the viscosity saturation as the analogous to the thermal conduction saturation (Cowie & McKee, 1977), following the implementation done in Su et al. (2017):
@@
Here, $T_{I}$ is the ion temperature, $n_{I}$ the ion number density, $e$ the electric charge and $\ln{\Lambda}=37.8$ the Coulomb logarithm. This way we make sure that when $l_{v}<\lambda_{I}$ (i.e. $|\sigma_{ij}|\gg 1$), the shear viscosity coefficient saturates, avoiding unphysical large values of the viscous stress tensor $\sigma$.
@@ 2.3 Simulation Setup
We also include magnetic fields based on the implementation of Bonafede et al. (2011) and Stasyszyn et al. (2013), with an initial seed of $B_{\mathrm{seed}}=10^{-12}$ G (comoving) in the $x$-direction, which corresponds to $B_{\mathrm{ini,ph}}=B_{\mathrm{seed}}\cdot(1+z_{\mathrm{ini}})^{2}=1.98\times 10^{-8}$ G in physical units. The choice of this initial magnetic field leads to a saturation of the dynamo at $\sim$1.5 at this resolution (Steinwandel et al., 2022, Marin-Gilabert et al, in prep). Our simulations also include anisotropic thermal conductivity via a bi-conjugate gradient solver (Arth et al., 2017; Steinwandel et al., 2022).
@@ 2.4 Initial Conditions
We chose the g55, g16 and g63 clusters because they have a $M_{\mathrm{Vir}}$ [^3] The suffix Vir indicates the virialized gas, i.e. the gravitationally bound gas. $>10^{15}$ at $z=0$; g55 and g16 are expected to be unrelaxed, while g63 is very relaxed. To classify the clusters as relaxed or unrelaxed, we used the centre-of-mass offset, where we measured the separation between the density peak position and the center of mass within $R_{\mathrm{200c}}$ (Power et al., 2012; Cui et al., 2016). The adopted cosmological parameters are $\Omega_{\mathrm{0}}=0.24$, $\Omega_{\mathrm{\Lambda}}=0.76$, $\Omega_{\mathrm{b}}=0.04$, $h=0.72$ and $\sigma_{8}=0.8$. The initial redshift for all the simulations is $z_{\mathrm{ini}}=140$ and the mass resolution is $m_{\mathrm{gas}}=1.56\times 10^{7}$ in gas particles and $m_{\mathrm{dm}}=8.44\times 10^{7}$ in dark matter. The choice of this particle resolution is due to the convergence in the magnetic field amplification shown in Steinwandel et al. (2022) against a lower particle resolution. The details of each cluster at $z=0$ are shown in table 1.

### v2

@@ Abstract
The impact of viscosity in the Intracluster Medium (ICM) is still an open question in astrophysics. To address this problem, we have run a set of cosmological simulations of three galaxy clusters with a mass larger than $M_{\mathrm{Vir}}>10^{15}$ at $z=0$ using the SPMHD-code OpenGadget3. We aim to quantify the influence of viscosity and constrain its value in the ICM. Our results show significant morphological differences at small scales, temperature variations, and density fluctuations induced by viscosity. We observe a suppression of instabilities at small scales, resulting in a more filamentary structure and a larger amount of small structures due to the lack of mixing with the medium. The conversion of kinetic to internal energy leads to an increase of the virial temperature of the cluster of $\sim$5- 10%, while the denser regions remain cold. The amplitude of density and velocity fluctuations are found to increase with viscosity. However, comparison with observational data indicates that the simulations, regardless of the viscosity, match the observed slope of the amplitude of density fluctuations, challenging the direct constraint of viscosity solely through density fluctuations. Furthermore, the ratio of density to velocity fluctuations remains close to 1 regardless of the amount of viscosity, in agreement with the theoretical expectations. Our results show for the first time in a cosmological simulation of a galaxy cluster the effect of viscosity in the ICM, a study that is currently missing in the literature.
@@ 1 Introduction
The Intracluster Medium (ICM) is the dominant baryonic component visible in galaxy clusters, filling the gravitational potential with hot ($T\sim 10^{7}-10^{8}$ K) ionised, X-ray emitting gas. This gas is continually perturbed by galaxy motions (e.g. Faltenbacher et al., 2005), mergers (e.g. ZuHone, 2011; Iapichino et al., 2017), AGN outflows (e.g. O’Neill et al., 2009; Gaspari, 2015) and accretion of gas along filaments (e.g. Kravtsov & Borgani, 2012; Vallés-Pérez et al., 2021a). These processes inject energy at large scales (low wavenumber $k$), which decays towards higher $k$ modes in a Kolmogorov-like cascade, introducing turbulence on a wide range of scales (e.g. Kolmogorov, 1941, 1962). The energy is then dissipated into heat, affecting small scale processes such as cosmic ray re-acceleration (e.g. Fujita et al., 2003; Brunetti & Lazarian, 2007), star formation (e.g. Federrath, 2016) or magnetic field amplification (e.g. Kazantsev, 1968; Kulsrud & Anderson, 1992). These complex mechanisms over cosmological periods of time produce very complicated scenarios in which the transport properties of the gas are fundamental.
@@
The properties of these density perturbations do not only reflect the driving mechanisms that trigger them, but also depend on the microphysics of the ICM, specifically, for thermal conduction (Ruszkowski et al., 2011) and viscosity (Zhuravleva et al., 2019; Kunz et al., 2022). Zhuravleva et al. (2019) suggested to use observations of density fluctuations to constrain the amount of viscosity in galaxy clusters. The idea is to compare the amplitude of density fluctuations measured in observations against simulations with different amounts of viscosity. Using Coma observations and simulations of incompressible hydrodynamic turbulence carried out using Direct Numerical Simulations (DNS), Zhuravleva et al. (2019) concluded that the effective isotropic viscosity must be suppressed by at least a factor of $\sim$10 to $\sim$1000 with respect to the Spitzer value. This suppression could be the effect of magnetic fields, which reduce the effect of viscosity (ZuHone et al., 2015; Berlok et al., 2019; Squire et al., 2023), as well as an enhanced scattering rate due to plasma instabilities (Schekochihin & Cowley, 2006; Kunz et al., 2014; Berlok et al., 2021).
@@
The physics of the ICM has been studied extensively using simulations of isolated systems with more idealised setups (e.g. ZuHone et al., 2009; Brüggen et al., 2012; Mohapatra et al., 2020; Zhang et al., 2020) and cosmological simulations (e.g. Dolag et al., 2005; Teyssier et al., 2011; Vazza et al., 2016; Pakmor et al., 2023; Steinwandel et al., 2024). Similarly, viscosity in the context of ICM has also been extensively studied with simulations, focusing on its effect on AGN-powered bubbles (Scannapieco & Brüggen, 2008; Dong & Stone, 2009), on the suppression of instabilities in cold fronts (Roediger & ZuHone, 2011; Suzuki et al., 2013; ZuHone et al., 2015) and on its impact on stripped galaxies (Roediger et al., 2015; Kraft et al., 2017).
@@
However, viscosity in the intracluster medium (ICM) has not been extensively studied using cosmological simulations. The only study in the literature, by Sijacki & Springel (2006), did not provide a detailed analysis of these effects. In our study, we conduct a set of cosmological simulations focusing on how viscosity affects massive clusters in a realistic scenario. We use different values of the isotropic Spitzer viscosity (Spitzer, 1962; Braginskii, 1965) and compare it with the non-viscous case. Our goal is to provide a general view of how the morphology of the cluster is affected by viscosity and how the virial temperature depends on the amount of viscosity. We also want to test the criteria to constrain viscosity suggested by Zhuravleva et al. (2019) and if this criterion holds in the context of the complex, realistic cosmological simulation. To this end, we compare our results with observations of density fluctuations from Heinrich et al. (2024). Additionally, we want to verify if the density fluctuations scale linearly with the velocity fluctuations, independently of the how viscous the ICM is. Although full Spitzer viscosity is larger than expected, due to the different mechanisms that suppress the viscous stress, we want to show this extreme case for a better understanding of the ongoing mechanisms. In a follow up work, we will show additional consequences of a viscous ICM such as changes of the magnetic field amplification, effects in the merger history or the different energy distribution, together with the time evolution of gas properties.
@@ 2.1 Theoretical considerations
$\rho\frac{\mathrm{d}\mathbf{v}}{\mathrm{d}t}+\nabla P=-\rho\nabla\Phi-\nabla\cdot\boldsymbol{\Pi}+\frac{(\nabla\times\mathbf{B})\times\mathbf{B}}{4\pi},$ (2)
$\frac{\mathrm{d}E}{\mathrm{d}t}+\mathbf{v}\cdot\nabla P+(E+P)\nabla\cdot\mathbf{v}-\nabla\cdot\frac{\mathbf{B}(\mathbf{v}\cdot\mathbf{B})}{4\pi}=\\ -\rho\mathbf{v}\cdot\nabla\Phi-\nabla\cdot\mathbf{Q}-\nabla\cdot\left(\boldsymbol{\Pi}\cdot\mathbf{v}\right)$ (3)
@@
is the Lagrangian derivative. $\rho$ is the gas density, $\mathbf{v}$ the velocity of the fluid, $\mathbf{B}$ the magnetic field, $\mathbf{Q}$ is the heat flux and $\Phi$ is the gravitational potential. $E$ is the energy per unit volume (kinetic + thermal + magnetic)
@@
with an adiabatic index of $\gamma=5/3$ for monoatomic gases. $\boldsymbol{\Pi}$ is the viscous stress tensor, which defines the behaviour of Spitzer viscosity as
$\boldsymbol{\Pi}=\eta\,\sigma_{ij}=\eta\,\left(\frac{\partial v_{i}}{\partial x_{j}}+\frac{\partial v_{j}}{\partial x_{i}}-\frac{2}{3}\delta_{ij}\frac{\partial v_{k}}{\partial x_{k}}\right)\,,$ (8)
where we have dropped the bulk viscosity term, since it is related to the degrees of freedom of molecular rotations, being zero for monoatomic gases (see e.g Zeldovich & Raizer, 1967; Pitaevskii & Lifshitz, 1981). The viscous stress tensor is equal to the rate of strain tensor ($\sigma_{ij}$) multiplied by the shear viscosity coefficient $\eta$, defined as
@@
Assuming subsonic gas motions within the galaxy cluster ($v\ll c_{s}$) and a stratified atmosphere, density perturbations are expected to be proportional to the one-component velocity at each scale $l=1/k$
$\left(\frac{\delta\rho}{\rho}\right)_{k}^{2}=\eta_{\rho}^{2}\left(\frac{v_{1\mathrm{D}}}{c_{s}}\right)_{k}^{2}=\eta_{\rho}^{2}\,\mathcal{M}_{\mathrm{1D},k}^{2}\,.$ (10)
$c_{s}$ is the soundspeed of the medium and $\mathcal{M}_{\mathrm{1D},k}$ the 1D mach number. A complete derivation can be found in Zhuravleva et al. (2014).
@@ 2.2 Viscosity saturation
A proper treatment of viscosity requires the implementation of a viscosity saturation to avoid unphysical results. The viscous stress saturates when the scale over which the velocity varies becomes smaller than the ion mean free path. This means that the viscous momentum transfer propagates faster than the mean soundspeed of the medium, overestimating this momentum transfer (Sarazin, 1986).
To avoid this, we need to introduce a viscosity saturation that limits the momentum propagation and ensures a smooth transition from the non-saturated to the saturated state. We do this in such a way that the momentum transfer of the saturated viscosity propagates at a velocity comparable to the soundspeed of the medium. We define the velocity length scale as $l_{v}=2\,c_{s}/|\sigma_{ij}|$, where
$|\sigma_{ij}|=\sqrt{\mathrm{tr}(\sigma_{ij}^{2})}=\sqrt{\mathrm{tr}(\sigma_{ij}\cdot\sigma_{ij})}=\\ \sqrt{\mathrm{tr}(\sigma_{xx}^{2}+\sigma_{yy}^{2}+\sigma_{zz}^{2}+2\sigma_{xy}^{2}+2\sigma_{xz}^{2}+2\sigma_{yz}^{2})}$ (11)
is the strength of the rate of strain tensor of the viscous stress tensor. We now introduce the viscosity saturation as the analogous to the thermal conduction saturation (Cowie & McKee, 1977), following the implementation done in Su et al. (2017):
@@
Here, $T_{I}$ is the ion temperature, $n_{I}$ the ion number density, $e$ the electric charge and $\ln{\Lambda}=37.8$ the Coulomb logarithm. This way we make sure that when $l_{v}<\lambda_{I}$, the shear viscosity coefficient saturates, avoiding unphysical large values of the viscous stress tensor $\boldsymbol{\Pi}$ and numerical issues like extremely small timesteps.
@@ 2.3 Simulation Setup
We also include magnetic fields based on the implementation of Bonafede et al. (2011) and Stasyszyn et al. (2013), with an initial seed of $B_{\mathrm{seed}}=10^{-12}$ G (comoving) in the $x$-direction, which corresponds to an initial magnetic field of $B_{\mathrm{ini,ph}}=B_{\mathrm{seed}}\cdot(1+z_{\mathrm{ini}})^{2}=1.98\times 10^{-8}$ G in physical units. The choice of this initial magnetic field leads to a saturation of the dynamo at $\sim$1.5 at this resolution (Steinwandel et al., 2022, Marin-Gilabert et al, in prep). Our simulations also include anisotropic thermal conductivity via a bi-conjugate gradient solver (Arth et al., 2017; Steinwandel et al., 2022).
@@ 2.4 Initial Conditions
We chose the g55, g16 and g63 clusters because they have a $M_{\mathrm{Vir}}$ [^2] The suffix Vir indicates the virialized gas, i.e. the gas within the galaxy cluster that satisfies the virial theorem. $>10^{15}$ at $z=0$; g55 and g16 are expected to be unrelaxed, while g63 is very relaxed. To classify the clusters as relaxed or unrelaxed, we used the centre-of-mass offset, where we measured the separation between the density peak position and the center of mass within $R_{\mathrm{200c}}$ (Power et al., 2012; Cui et al., 2016). The adopted cosmological parameters are $\Omega_{\mathrm{0}}=0.24$, $\Omega_{\mathrm{\Lambda}}=0.76$, $\Omega_{\mathrm{b}}=0.04$, $h=0.72$ and $\sigma_{8}=0.8$. The initial redshift for all the simulations is $z_{\mathrm{ini}}=140$ and the mass resolution is $m_{\mathrm{gas}}=1.56\times 10^{7}$ in gas particles and $m_{\mathrm{dm}}=8.44\times 10^{7}$ in dark matter. The choice of this particle resolution is due to the convergence in the magnetic field amplification shown in Steinwandel et al. (2022) against a lower particle resolution. The details of each cluster at $z=0$ are shown in table 1.